OpenStudy (anonymous):

If (2 − 3i) + (x + yi) = 6, what is x + yi? 4 + 3i 4 − 3i -4 − 3i -4 + 3i 4x + 3i

OpenStudy (anonymous):

i solved it and i didn't get nun of your answer choice

OpenStudy (anonymous):

this is what i got Let's solve for i. 2−3i+x+yi=6 Step 1: Add -2 to both sides. iy−3i+x+2+−2=6+−2 iy−3i+x=4 Step 2: Add -x to both sides. iy−3i+x+−x=4+−x iy−3i=−x+4 Step 3: Factor out variable i. i(y−3)=−x+4 Step 4: Divide both sides by y-3. i(y−3)y−3=−x+4y−3 i=−x+4y−3 Answer: i=−x+4y−3

OpenStudy (anonymous):

so i would probably say B

OpenStudy (anonymous):

okay i see how to do these now thx